# Similarity-Distance-Magnitude Activations

### v2

@@ 4 Experiments
We provide controlled comparisons of our proposed methods over representative LMs and tasks, systematically ablating relevant components, holding the data and underlying LM constant, ceteris paribus. We consider in-distribution, co-variate shifted, and out-of-distribution test sets. We consider representative estimators over the existing LM architecture (i.e., without additional parameters); with CNN adaptors; and with the $\operatorname{\text{{sdm}}}$ activation layer.
@@ 4.1 Task: Binary Sentiment Classification
Our first task is predicting the sentiment of movie reviews. We use the commonly used benchmark data of Maas et al. (2011). This is a binary classification task with $y\in\{0={\rm{negative}},1={\rm{positive}}\}$. $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ are constructed from a total of 18k instances. The held-out set for evaluation, $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=1583$, is from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. This is a well-studied task for which the surface-level signals correlated with the target labels are expected to be effectively modeled by large parameter LMs; as such, relatively high task accuracies are expected. We use the label $\operatorname{\text{{Sentiment}}}$ for the in-distribution test set.
@@
$\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$.
In the Appendix, we consider two additional out-of-distribution challenge test sets, $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, constructed by randomly shuffling the input documents for each of $\operatorname{\text{{Sentiment}}}$ and $\operatorname{\text{{SentimentOOD}}}$, respectively. The semantics of the original labels are maintained. This requires the models and estimators to attempt a sentiment classification over the bag-of-words input, or reject the classification. This represents the setting where an LM is given far out-of-distribution input, and additionally provides a control on test-set contamination of the underlying LMs, which due to the shuffling, are relatively unlikely to have seen all of the long, contiguous n-gram sequences from these documents in training or fine-tuning.
Prompt.
For this task, we prompt the LMs for a binary classification (Yes or No) as follows: Here is a movie review. <review> DOCUMENT </review> Is the sentiment of the movie review positive? Answer Yes if the sentiment is positive. Answer No if the sentiment is negative. Start your response with Yes or No.
We replace DOCUMENT with the corresponding text for each instance.
@@ 4.2 Task: $\operatorname{\text{{Factcheck}}}$
As a more challenging binary classification task for LMs, we consider the fact check data of Azaria & Mitchell (2023). The training and calibration sets, a combined total of 6k instances, consist of single sentence statements that have been semi-automatically generated via templates and a knowledge base. The task is to determine whether the statement is true or false, $y\in\{0={\rm{false}},1={\rm{true}}\}$. The held-out eval set ($\operatorname{\text{{Factcheck}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=245$, the focus of our analysis, has been constructed by having an LM generate a statement continued from a true statement not otherwise in the dataset. These evaluation statements are checked manually and assigned labels by human annotators. In addition to being a relatively challenging task that evaluates—at least in principle—the latent knowledge stored within an LM’s parameters, the test set is representative of the types of co-variate shifts over high-dimensional inputs that can be problematic for real applications, and challenging to characterize without model assistance and ground-truth labels. It was observed in Azaria & Mitchell (2023) that the accuracy of existing LM classifiers is lower on this generated, held-out test set compared to the calibration set. However, these test sentences would seem to also be simple true-false statements, reflecting that it is not necessarily straightforward for a human user to detect distribution shifts over high-dimensional inputs. As such, we seek for our models and estimators to reflect such shifts via the predictive uncertainty.
@@
As with the sentiment task, in the Appendix we also consider an additional out-of-distribution challenge test set, $\operatorname{\text{{FactcheckShuffled}}}$, constructed by randomly shuffling the input documents of $\operatorname{\text{{Factcheck}}}$. Since the task prompt (below) seeks to classify errors, we set the ground-truth labels of the shuffled counterparts to $y=0$.
Prompt.
Similar to the sentiment task, we prompt the LMs for a binary classification (Yes or No) as follows: Here is a statement that may contain errors. <statement> DOCUMENT </statement> Is the statement true? Answer Yes if the statement is true. Answer No if the statement is false. Start your response with Yes or No.
We replace DOCUMENT with the corresponding text for each instance.
@@ 4.3 Models
We consider two representative, publicly-available decoder-only Transformer-based language models: The 3.8 billion-parameter Phi-3.5-mini-instruct model ($\operatorname{\text{{phi3.5}}}$) (Abdin et al., 2024), and the 47 billion-parameter Mixtral 8x7B Instruct v0.1 mixture-of-experts model ($\operatorname{\text{{Mixtral8x7B}}}$) (Jiang et al., 2024). The parameters of these models stay fixed in all experiments.
@@
For both models, we take as $\bm{h}$} the concatenation of the final-layer hidden state of the final sequence position (i.e., the hidden state that is the input to the linear-layer over the output vocabulary for the Yes or No generation) with the mean over all final-layer hidden states. For $\operatorname{\text{{phi3.5}}}$, this results in ${\bm{h}}\in\mathbb{R}^{6144}$, and for $\operatorname{\text{{Mixtral8x7B}}}$, ${\bm{h}}\in\mathbb{R}^{8192}$.
@@ 4.4 Estimators
Holding each underlying LM constant, we examine representative classes of estimators used with neural networks, setting $\alpha=0.95$ for all experiments. At the most basic, but also perhaps the most commonly used in practice, representing the absence of a post-hoc calibration method, we simply threshold the output, $\operatorname{softmax}({\bm{z}})\geq\alpha$, where the temperature $\tau=1$. As an established empirical approach for calibrating neural networks, we provide a comparison to temperature scaling (Guo et al., 2017), a single parameter version of post-hoc Platt-scaling (Platt, 1999), with the label $\operatorname{\text{{tempScaling}}}$. In this case, the estimator is the thresholding of the output $\operatorname{softmax}({\bm{z}};\tau)\geq\alpha$ after learning a value for $\tau$ over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We also provide a comparison to two representative conformal predictors, the $\operatorname{\textsc{APS}}$ method of Romano et al. (2020) and the adaptiveness-optimized $\operatorname{\textsc{RAPS}}$ algorithm of Angelopoulos et al. (2021). The admission criteria for the $\operatorname{\textsc{APS}}$ and $\operatorname{\textsc{RAPS}}$ estimators is prediction sets of size 1, at the $0.05$ level (i.e., $1-\alpha$, as defined here). We consider these estimators over the logits corresponding to the Yes and No indexes of the output linear-layer of the underlying LM ($\operatorname{\text{{phi3.5}}}$ and $\operatorname{\text{{Mixtral8x7B}}}$), which provides a reference point without introducing additional adaptor layers. We also consider these baselines over 1-D CNN adaptors over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+adaptor}}}$ and $\operatorname{\text{{Mixtral8x7B+adaptor}}}$). These baseline adaptors are identical to those used in the corresponding $\operatorname{\text{{sdm}}}$ activation layers, with $M=1000$, and are similarly trained for $J=10$ iterations of 200 epochs, but unlike the $\operatorname{\text{{sdm}}}$ activation layers, the stopping criteria is the minimum balanced (across classes) average cross-entropy loss.
@@
We then compare to the final-layer $\operatorname{\text{{sdm}}}$ activations over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+sdm}}}$ and $\operatorname{\text{{Mixtral8x7B+sdm}}}$). For reference, we provide the result of thresholding a $\operatorname{softmax}$ over these adaptors at $\alpha$, as above, as well as a thresholded $\operatorname{softmax}$ that simply treats $\operatorname{\text{$d$}}$ as the inverse-temperature, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$, which is equivalent to setting q=e2q=e-2 in the $\operatorname{\text{{sdm}}}$ activation. We also consider an analogous threshold over the activation output, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})\geq\alpha$, for which we use $\operatorname{\text{{sdm}}}$ as the estimator label. Finally, we use the label $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ for the selection criteria proposed in Eq. 9.

### v3

@@ 4 Experiments
We provide controlled comparisons of our proposed methods over representative LMs and tasks, systematically ablating relevant components, holding the data and underlying LM constant, ceteris paribus. We consider in-distribution, co-variate shifted, and out-of-distribution test sets. We consider representative estimators over the existing LM architecture (i.e., without additional parameters); with CNN adaptors; and with the $\operatorname{\text{{sdm}}}$ activation layer. Additional details are provided in the Appendix.
@@ 4.1 Task: Binary Sentiment Classification
Our first task is predicting the sentiment of movie reviews. We use the commonly used benchmark data of Maas et al. (2011). This is a binary classification task with $y\in\{0={\rm{negative}},1={\rm{positive}}\}$. $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$ are constructed from a total of 18k instances. The held-out set for evaluation ($\operatorname{\text{{Sentiment}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=1583$, is from the same distribution as $\operatorname{\text{${\mathcal{D}}_{\rm{tr}}$}}$ and $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$.
@@
Far OOD Challenge Sets.
In Appendix A.4, we consider two additional out-of-distribution challenge test sets, $\operatorname{\text{{SentimentShuffled}}}$ and $\operatorname{\text{{SentimentOODShuffled}}}$, constructed by randomly shuffling the input documents for each of $\operatorname{\text{{Sentiment}}}$ and $\operatorname{\text{{SentimentOOD}}}$, respectively.
@@ 4.2 Task: $\operatorname{\text{{Factcheck}}}$
As a more challenging binary classification task for LMs, we consider the fact check data of Azaria and Mitchell (2023). The training and calibration sets, a combined total of 6k instances, consist of single sentence statements that have been semi-automatically generated via templates and a knowledge base. The task is to determine whether the statement is true or false, $y\in\{0={\rm{false}},1={\rm{true}}\}$. The held-out eval set ($\operatorname{\text{{Factcheck}}}$), $|\operatorname{\text{${\mathcal{D}}_{\rm{te}}$}}|=245$, the focus of our analysis, has been constructed by having an LM generate a statement continued from a true statement not otherwise in the dataset. These evaluation statements are checked manually and assigned labels by human annotators. In addition to being a relatively challenging task that evaluates—at least in principle—the latent knowledge stored within an LM’s parameters, the test set is representative of the types of co-variate shifts over high-dimensional inputs that can be problematic for real applications, and challenging to characterize without model assistance and ground-truth labels. It was observed in Azaria and Mitchell (2023) that the accuracy of existing LM classifiers is lower on this generated, held-out test set compared to the calibration set. However, these test sentences would seem to also be simple true-false statements, reflecting that it is not necessarily straightforward for a human user to detect distribution shifts over high-dimensional inputs. As such, we seek for our models and estimators to reflect such shifts via the predictive uncertainty.
@@
As with the sentiment task, in Appendix A.4 we also consider an additional out-of-distribution challenge test set, $\operatorname{\text{{FactcheckShuffled}}}$, constructed by randomly shuffling the input documents of $\operatorname{\text{{Factcheck}}}$.
@@ 4.3 Models
We consider two representative, open-weight decoder-only Transformer-based LMs, the parameters of which stay fixed: The 3.8 billion-parameter Phi-3.5-mini-instruct model ($\operatorname{\text{{phi3.5}}}$) (Abdin et al., 2024), and the 47 billion-parameter Mixtral 8x7B Instruct v0.1 mixture-of-experts model ($\operatorname{\text{{Mixtral8x7B}}}$) (Jiang et al., 2024).
@@
We take as $\bm{h}$} the concatenation of the final-layer hidden state of the final sequence position (i.e., the hidden state that is the input to the linear-layer over the output vocabulary for the Yes or No generation) with the mean over all final-layer hidden states. For $\operatorname{\text{{phi3.5}}}$, this results in ${\bm{h}}\in\mathbb{R}^{6144}$, and for $\operatorname{\text{{Mixtral8x7B}}}$, ${\bm{h}}\in\mathbb{R}^{8192}$.
@@ 4.4 Estimators
We examine representative Frequentist, Bayesian, and empirically-motivated classes of estimators used with neural networks, setting $\alpha=0.95$ for all experiments. At the most basic, but perhaps the most commonly used in practice, representing the absence of a post-hoc calibration method, we simply threshold the output, $\operatorname{softmax}({\bm{z}})\geq\alpha$, where the temperature $\tau=1$. For this, we use the label $\operatorname{softmax}$. As an established empirical approach for calibrating neural networks, we provide a comparison to temperature scaling (Guo et al., 2017), a single parameter version of post-hoc Platt-scaling (Platt, 1999), with the label $\operatorname{\text{{tempScaling}}}$. In this case, the estimator is the thresholding of the output $\operatorname{softmax}({\bm{z}};\tau)\geq\alpha$ after learning a value for $\tau$ over $\operatorname{\text{${\mathcal{D}}_{\rm{ca}}$}}$. We also provide a comparison to two representative conformal predictors, the $\operatorname{\textsc{APS}}$ method of Romano et al. (2020) and the adaptiveness-optimized $\operatorname{\textsc{RAPS}}$ algorithm of Angelopoulos et al. (2021). The admission criteria for the $\operatorname{\textsc{APS}}$ and $\operatorname{\textsc{RAPS}}$ estimators is prediction sets of size 1, at the $0.05$ level (i.e., $1-\alpha$, as defined here). We consider these estimators over the logits corresponding to the Yes and No indexes of the output linear-layer of the underlying LM ($\operatorname{\text{{phi3.5}}}$ and $\operatorname{\text{{Mixtral8x7B}}}$), which provides a reference point without introducing additional adaptor layers. We also consider these baselines over 1-D CNN adaptors over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+adaptor}}}$ and $\operatorname{\text{{Mixtral8x7B+adaptor}}}$). These baseline adaptors are identical to those used in the corresponding $\operatorname{\text{{sdm}}}$ activation layers, with $M=1000$, and are similarly trained for $J=10$ iterations of 200 epochs, but unlike the $\operatorname{\text{{sdm}}}$ activation layers, the stopping criteria is the minimum balanced (across classes) average cross-entropy loss.
We further compare to variational Bayesian last-layer neural networks ($\operatorname{\text{{vbll}}}$), a computationally efficient Bayesian approach (Harrison et al., 2024). We consider the discriminative versions over Phi-3.5-mini-instruct and Mixtral 8x7B Instruct v0.1 ($\operatorname{\text{{phi3.5+DiscVBLLMLP}}}$ and $\operatorname{\text{{Mixtral8x7B+DiscVBLLMLP}}}$, respectively) in the main text, with additional comparisons in Appendix A.7.
@@
We then compare to the final-layer $\operatorname{\text{{sdm}}}$ activations over $\bm{h}$} of each LM ($\operatorname{\text{{phi3.5+sdm}}}$ and $\operatorname{\text{{Mixtral8x7B+sdm}}}$). For reference, we provide the result of thresholding a $\operatorname{softmax}$ over these adaptors at $\alpha$, as above, as well as a thresholded $\operatorname{softmax}$ that simply treats $\operatorname{\text{$d$}}$ as the inverse-temperature, $\operatorname{\text{{softmax}}}(\operatorname{\text{$d$}}\cdot{\bm{z}}^{\prime})$, which is equivalent to setting q=e2q=e-2 in the $\operatorname{\text{{sdm}}}$ activation. We consider an analogous threshold over the activation output, $\operatorname{\text{{sdm}}}({\bm{z}}^{\prime})\geq\alpha$, for which we use $\operatorname{\text{{sdm}}}_{\text{$\alpha$}}$ as the estimator label. Finally, for the eponymous $\operatorname{\text{{sdm}}}$ estimator” using the selection criteria of Eq. 9, we use the label $\operatorname{\text{{sdm}}}_{\operatorname{\text{{HR}}}}$ in the results tables.